# LexGenius: An Expert-Level Benchmark for Large Language Models in Legal General Intelligence

### v1

@@ Abstract
Figure 1: Comparison of LLM and human legal experts shows that humans outperform LLMs in all dimensions.
@@ 1 Introduction
“Laws are the expression of the will of the people.”
— Montesquieu
@@
Legal general intelligence is the capacity of general AI to perform with expert-level ability across complex legal contexts (e.g., hard tasks, soft intelligence) (Nasir et al., 2024; Kant et al., 2025; Zhou et al., 2025). It involves the precise interpretation of legal provisions, sound inference based on complex factual scenarios Zhang et al. (2025b); Li et al. (2025e); Shen et al. (2025), the resolution of conflicts among rules from multiple interrelated legal domains, and the ability to make normatively binding judgments Yue et al. (2024a); Zhang et al. (2025d); Luo et al. (2025b) in uncertain and ethically sensitive contexts Kim et al. (2025); Huang et al. (2023); Liu et al. (2025c). Legal GI is not just whether AI knows the law, but whether it can participate in the normative structure of legal systems, opening the door to its integration into legal order.
@@
In recent years, LLMs have demonstrated strong performance across general language tasks Mao et al. (2024); Zheng et al. (2025). This progress has catalyzed a surge of interest in adapting LLMs to the legal domain, aiming to tackle challenges, such as legal question-answering  Su et al. (2025); Fei et al. (2024), case analysis Zhang et al. (2025a); Li et al. (2025a), and judgment prediction Liu et al. (2025b); Xie et al. (2025). To assess the legal reasoning capabilities of LLMs, several benchmarks, such as LegalBench Guha et al. (2023), LexEval Li et al. (2024a), and LexGLUE Chalkidis et al. (2021); Jia et al. (2025), are introduced. These benchmarks provide a critical foundation for evaluating, improving, and advancing the legal capabilities of large language models  Luo et al. (2025a, c).
@@
However, the existing benchmarks encounter the following limitations: (1) Legal intelligence has not yet entered the second half of AI. Current benchmarks Chalkidis et al. (2021); Fei et al. (2024) focus on technical tasks while neglecting soft legal intelligence, such as ethical judgment, the law–morality boundary, and societal impact assessment Wang et al. (2024); Cambria et al. (2024). (2) Data contamination and lack of comprehensiveness. Static, publicly available legal benchmarks risk data leakage (Wu et al., 2025) and fail to assess models on dynamic reasoning or novel legal scenarios, leading to overstated and unreliable evaluations. (3) Lack of a structured framework for comprehensively assessing legal intelligence abilities. Outcome-focused benchmarks overlook legal reasoning stages, blurring the line between true understanding and pattern mimicry Cui et al. (2023); Hassani et al. (2025); Chen et al. (2024).
@@
To address the above limitations, we propose LexGenius, a comprehensive benchmark to assess legal general intelligence for LLMs. First, we rethink the evaluation of legal intelligence for LLMs Thakur et al. (2024). Recognizing that existing benchmarks Li et al. (2024b); Wang et al. (2024); Chang et al. (2024); Xu et al. (2025a); Fei et al. (2024) overlook aspects of legal soft intelligence, our framework explicitly incorporates tests of capabilities such as ethical judgment, moral-legal boundaries, and social impact. We have developed a new collection of 8,385 standardized legal multiple-choice questions (MCQs), covering civil, criminal, and commercial law. To ensure legal accuracy, all questions and answers are refined through professional review. These MCQs assess multi-dimensional competencies Cambria et al. (2024); Corfmat et al. (2025) relevant to legal intelligence. Furthermore, to address cognitive coverage limitations in existing benchmarks, we propose a framework structured around the dimensions of legal theory and practice, organizing tasks and abilities to reflect real-world legal intelligence Gursoy and Cai (2025). Focusing on Chinese laws ensures a robust and meaningful assessment, as distinct legal systems would otherwise dilute the evaluation.
@@
Leveraging LexGenius, we evaluated 12 state-of-the-art (SOTA) LLMs and 2 prompting strategies, including naive and chain-of-thought (CoT) prompting Kojima et al. (2022). A baseline performance, constructed by 6 legal professionals, was also established for comparison. Results show that even the top-performing LLM, DeepSeek-R1 Bi et al. (2024), exhibits a significant gap compared to human experts across various legal general intelligence abilities Yao et al. (2025b); Hannah et al. (2025); Dong et al. (2025), as shown in Figure 1. In summary, our contributions in this work include:
@@ 2 Related Work
Legal Benchmarks. Recently, a series of legal benchmarks have emerged (see Table 1). They have made significant contributions to evaluating LLMs’ performance Kanapala et al. (2019); Yao et al. (2025a), including retrieval (STARD, LeCaRD) Su et al. (2024); Li et al. (2024c), question answering (JEC-QA, Legal CQA) Zhong et al. (2020); Askari et al. (2022), classification (LexGLUE) Chalkidis et al. (2021), reasoning (LegalBench, LexEval) Guha et al. (2023); Li et al. (2024a), and others (Laiw and LawBench) Dai et al. (2025); Fei et al. (2024). However, most benchmarks remain task-oriented and outcome-focused, offering limited insight into the underlying legal general intelligence of LLMs  Yue et al. (2024b).
Expert-level Benchmarks. To usher in the second half of AI, a series of expert-level benchmarks for evaluating LLMs have emerged across various domains Cao et al. (2025); Ni et al. (2025); Li et al. (2025c, d): PhysBench Chow et al. (2025) and PhysReason Zhang et al. (2025c) enhance LLMs’ understanding of physics; MedXpertQA Zuo et al. (2025) and MedagentsbenchTang et al. (2025) focus on medical knowledge; UGMathBench Xu et al. (2025b) assesses math reasoning; and UniToMBench Thiyagarajan et al. (2025) improves theory of mind. Benchmarks like ShotBench Liu et al. (2025a), FinTMMBench Zhu et al. (2025), and Chengyu-Bench Fu et al. (2025) evaluate other fields. However, an expert-level benchmark for legal intelligence is absent.  Wang et al. (2025).
@@ 3 LexGenius Framework
3.1 Dimension: Education and Career Drive
The seven legal dimensions of LexGenius are based on Bloom’s Taxonomy of Educational Objectives Bloom et al. (1956), covering the cognitive hierarchy of remembering, understanding, applying, analyzing, evaluating, and creating, alongside the modular model used in legal evaluations across countries, focusing on normative understanding, rule application, procedural operation, and value judgment Wu and Chan (2012); Moon (2020); Parsons et al. (2024). In the hierarchy, remembering and understanding correspond to legal understanding, applying to legal application, analyzing to legal reasoning, evaluating to legal ethics and law and society, creating to advanced arguments, legal language to clarity, and judicial practice to procedural integrity, forming a framework aligned with cognitive principles and professional needs.
@@ 3.1 Dimension: Education and Career Drive
3.2 Task: Theory and Practice Coexist
Further, based on Legal Hermeneutics Leyh (2021) and Problem-Solving Cycle theory Stein (1993), we decompose LexGenius’s 7 legal intelligence dimensions into 11 tasks. These tasks align with common legal practice requirements and focus on textual deconstruction, case adaptation, and procedural implementation. Legal Hermeneutics guides the understanding of provisions, critical analysis of texts, and terminology, ensuring accuracy in interpretation. Problem-Solving Cycle theory simulates legal practice, driving reasoning and application analysis for problem-solving, legal and ethical judgment, moral boundaries for value calibration, case reasoning, judicial procedure understanding for validation, and legal impact and social change review, forming a task system for legal problem-solving.
@@ 3.3 Ability: Constructivist Learning-based
Furthermore, based on Constructivist Learning Theory Ariati et al. (2025), we extract twenty atomic legal intelligence abilities from the eleven tasks. The theory shifts from outcome assessment to capturing knowledge paths through cognitive traces, simulating real legal scenarios to ensure that the evaluation reflects true professional abilities while aligning cognitive principles with occupational needs. The hierarchy from dimensions to abilities allows LexGenius to perform a detailed, multi-dimensional assessment of LLMs’ legal intelligence, supporting evaluation and optimization.
@@ 4.1 Construction Principles
To avoid data leakage and contamination, LexGenius was built from scratch, using recent Chinese legal exam questions and judgment cases. We avoided reusing existing legal datasets to minimize contamination risks and ensure originality. A structured legal competency framework, developed by experienced legal experts, was used to comprehensively cover core legal intelligence abilities Li et al. (2025f). To ensure benchmark effectiveness, a structured review process was established Mohammadi et al. (2025); Li et al. (2025b). Reviewers are master’s candidates in law, systematically trained and thoroughly familiar with key regulations, case frameworks, and legal reasoning methods.
@@ 4.2 Construction Workflow
Step 2: Construction of MCQs. There are two methods for constructing multiple-choice questions, both based on LLM. One is to screen and modify legal examination questions. Questions meeting the defined legal abilities are retained; others are modified using LLM. For questions related to legal soft intelligence Schiff et al. (2025), LLM is used to generate them. The first method selects and modifies questions from the legal question bank: MCQs with a single correct answer are retained. In contrast, those with multiple answers are adapted using LLMs and prompt templates to ensure fairness and difficulty. For LLM-generated MCQs, we designed prompt templates with task constraints, ability descriptions, and examples to guide question generation based on legal cases, ensuring unique answers and a clear legal basis.
@@ 4.4 Evaluation Method
To evaluate the LLM’s legal intelligence, LexGenius adopts a three-level framework Kahan (2015); Huang et al. (2024). The dimension level categorizes tasks into legal cognition areas, providing insight into performance. The task level breaks dimensions into real-world tasks, testing the application, while the ability level evaluates legal abilities, identifying performance differences. At the ability level, scores $A_{i,j,k}$ are the average correctness of MCQs in each ability, where $A_{i,j,k}=\frac{1}{n_{i,j,k}}\sum_{m=1}^{n_{i,j,k}}C_{i,j,k,m}$, with $C_{i,j,k,m}$ the correctness of the $m$-th MCQ in the $k$-th ability of the $j$-th task in the $i$-th dimension, and $n_{i,j,k}$ the number of MCQs for that ability. At the task level, the task score $T_{i,j}$ is the average of ability scores within the task, calculated as $T_{i,j}=\frac{1}{m_{i,j}}\sum_{k=1}^{m_{i,j}}A_{i,j,k}$, where $m_{i,j}$ is the number of abilities in the $j$-th task of the $i$-th dimension. At the dimension level, the dimension score $D_{i}$ is the average of task scores, expressed as $D_{i}=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}T_{i,j}$, where $n_{i}$ is the number of tasks in the $i$-th dimension.
@@ 5.1 Experimental Setup
We evaluated twelve SOTA LLMs with LexGenius, which include DeepSeek-LLM-7B-Chat (DeepSeek-7B) Bi et al. (2024), Qwen-2.5-7B-Instruct (Qwen-2.5-7B) Hui et al. (2024), Qwen-2.5-1.5B-Instruct (Qwen-2.5-1.5B) Hui et al. (2024), Qwen-3-8B Yang et al. (2025), Qwen-3-4B Yang et al. (2025), GLM-4-9B-Chat (GLM-4-9B) GLM et al. (2024), LLaMA-3.2-1B-Instruct (LLaMA-3.2-1B) Grattafiori et al. (2024), LLaMA-3.2-8B-Instruct (LLaMA-3.2-8B) Grattafiori et al. (2024), DeepSeek-R1 Guo et al. (2025), DeepSeek-V3 Liu et al. (2024), GPT-4o mini Hurst et al. (2024), and GPT‑4.1 nano Brown et al. (2020). We followed the official protocols, using official APIs or LLM weights where applicable. The evaluation utilized two types of prompts: the first type was the Naive prompt; the second was the CoT prompt, encouraging the LLMs to perform step-by-step reasoning. To prevent potential bias, we shuffled the answer options twice and averaged the scores of each LLM.
@@ 5.5 With Different Enhanced Methods (RQ4)
As shown in Table 5, the comparison results (more details are in Appendix E) reveal a triple decoupling phenomenon in LLM legal intelligence: (i) Scale-performance decoupling shows that the Qwen2.5 series outperforms the Qwen3 series, with legal intelligence exhibiting a non-linear relationship with model size, indicating that domain-specific pretraining, not scale, drives legal abilities. (ii) Reasoning paradigm decoupling shows that CoT leads to negative transfer in legal tasks due to deterministic constraints, causing probabilistic exploration and semantic drift in closed-solution tasks. (iii) Optimization strategy decoupling shows that supervised fine-tuning benefits weak LLMs but causes catastrophic forgetting in strong ones, while the RAG method reveals the orthogonality of knowledge retrieval and reasoning. Only GRPO (reinforcement learning) achieves stable improvement by aligning reward and evaluative capacity.
@@ A.1 Design motivation
Legal general intelligence is not a stack of tasks, but a simulation of cognitive collaborative chains. Existing evaluation frameworks for legal intelligence often operate at the level of classification or question-answering tasks, presenting only the model’s macro-level accuracy on isolated benchmarks. Such metrics fail to uncover where a model goes wrong within the complex cognitive chain of legal decision-making. Legal intelligence is not a mere sum of discrete tasks—it is an organic coordination of systematic cognitive capabilities, spanning multiple stages such as statutory interpretation, fact extraction, rule adaptation, ethical judgment, and social impact assessment. The performance of an LLM should therefore be decomposed into a multi-stage flow of abilities, rather than treated as a monolithic output. Accordingly, an effective evaluation framework must faithfully map this cognitive chain, avoiding the pitfall of reducing legal intelligence to mere task-solving while overlooking how the model actually thinks.
@@
From performance reporting to capability diagnosis and explanation. Traditional metrics such as accuracy or F1 score can indicate whether a model’s output is right or wrong, but they fail to address a more fundamental question: At which cognitive stage did the model fail? Was it a semantic misunderstanding? A misapplication of legal rules? Or an ethical blind spot? LexGenius introduces 20 atomic legal general intelligence abilities and establishes an interpretable, traceable, and auditable evaluation system through a tri-level mapping mechanism across the ability layer, task layer, and cognitive dimension. This structure not only reveals the model’s deficiencies in specific micro-level legal abilities but also provides an actionable feedback loop for subsequent capability-oriented training, prompt optimization, and safety enhancement.
@@
Enabling cross-stage cognitive analysis and transferability research. The complexity of legal reasoning lies in its chained structure: statutory semantics $\rightarrow$ case fact abstraction $\rightarrow$ rule application $\rightarrow$ ethical judgment $\rightarrow$ precedent alignment. An evaluation framework that fails to distinguish performance across these stages cannot effectively support research in multi-hop reasoning, chain-of-thought attention, or multi-task learning. Through hierarchical abstraction and modular decomposition, our framework standardizes this cognitive pathway, offering a clear experimental baseline for subsequent investigations into how models learn to transfer knowledge and whether they exhibit generalizable reasoning capabilities.
@@
Toward a professional-grade legal general intelligence evaluation paradigm. As LLMs begin to approach the professional thresholds of bar examinations and real-world legal practice, evaluation frameworks must likewise advance to a professional-grade level. LexGenius draws on standards from the National Legal Professional Qualification Examination and international bar exams, distinguishing among multiple specialized dimensions of legal competence, such as legal semantic understanding, norm alignment, ethical judgment, and procedural compliance. This framework breaks away from the general-purpose perspective of traditional NLP benchmarks, aligning instead with the practical demands of legal work. Such a professional, ability-oriented, structured evaluation not only reveals the current boundaries of model capabilities but also illuminates their potential trajectories of future development.

### v2

@@ Abstract
Figure 1: Comparison of the state-of-the-art LLMs and human legal experts illustrates that humans outperform LLMs in the seven legal intelligence dimensions.
@@ 1 Introduction
“The law is the expression of the general will.”
— Jean-Jacques Rousseau
@@
Legal general intelligence is the capacity of general AI to perform with expert-level ability across complex legal contexts (e.g., hard tasks, soft intelligence) (kant2025towards; zhou2025lawgpt). It involves the precise interpretation of legal provisions, sound inference based on complex factual scenarios zhang2025syler; li2025basis; SHEN2025102860, the resolution of conflicts among rules from multiple interrelated legal domains, and the ability to make normatively binding judgments yue2024circumstance; zhang2025rljp; luo2025hypergraphrag in uncertain and ethically sensitive contexts kim2025legisflow; huang2023lawyer; liu2025prompt. Legal general intelligence is not just whether AI knows the law, but whether it can participate in the normative structure of legal systems, thereby opening the door to its integration into legal order.
@@
In recent years, LLMs have demonstrated strong performance across general language tasks mao2024gpteval; zheng2025towards. This progress has catalyzed a surge of interest in adapting LLMs to the legal domain, aiming to tackle challenges, such as legal question-answering  su2025judge; fei2024lawbench, case analysis zhang2025citalaw; li2025unilr, and judgment prediction liu2025legal; xie2025lawchain. To assess the legal reasoning capabilities of LLMs, several benchmarks, such as LegalBench guha2023legalbench, LexEval li2024lexeval, and LexGLUE chalkidis2021lexglue; jia2025ready, are introduced. These benchmarks provide a critical foundation for evaluating, improving, and advancing the legal capabilities of large language models  luo2025graph; luo2025kbqao1.
@@
However, the existing benchmarks encounter the following limitations: (1) Legal intelligence has not yet entered the second half of AI. Current benchmarks chalkidis2021lexglue; fei2024lawbench focus on technical tasks while neglecting soft legal intelligence, such as ethical judgment, the law–morality boundary, and societal impact assessment wang2024legal; cambria2024xai. (2) Data contamination and lack of comprehensiveness. Static, publicly available legal benchmarks risk data leakage (wu2025antileak) and fail to assess models on dynamic reasoning or novel legal scenarios, leading to overstated and unreliable evaluations. (3) Lack of a structured framework for comprehensively assessing legal intelligence abilities. Outcome-focused benchmarks overlook legal reasoning stages, blurring the line between true understanding and pattern mimicry cui2023chatlaw; hassani2025empirical; chen2024survey.
@@
To address the above limitations, we propose LexGenius, a comprehensive benchmark to assess legal general intelligence for LLMs. First, we rethink the evaluation of legal intelligence for LLMs thakur2024judging. Recognizing that existing benchmarks li2024llms; wang2024legal; chang2024survey; xu2025towards; fei2024lawbench overlook aspects of legal soft intelligence, our framework explicitly incorporates tests of capabilities such as ethical judgment, moral-legal boundaries, and social impact. We have developed a new collection of 8,385 standardized legal multiple-choice questions (MCQs), covering civil, criminal, and commercial law. To ensure legal accuracy, all questions and answers are refined through professional review. These MCQs assess multi-dimensional competencies cambria2024xai; corfmat2025high relevant to legal intelligence. Furthermore, to address cognitive coverage limitations in existing benchmarks, we propose a framework structured around the dimensions of legal theory and practice, organizing tasks and abilities to reflect real-world legal intelligence gursoy2025artificial. Focusing on Chinese laws ensures a robust and meaningful assessment, as distinct legal systems would otherwise dilute the evaluation.
@@
Leveraging LexGenius, we evaluated 12 state-of-the-art (SOTA) LLMs and 2 prompting strategies, including naive and chain-of-thought (CoT) prompting kojima2022large. A baseline performance, constructed by 6 legal professionals, was also established for comparison. Results show that even the top-performing LLM, DeepSeek-R1 bi2024deepseek, exhibits a significant gap compared to human experts across various legal general intelligence abilities yao2025intelligent; hannah2025legal; dong2025safeguarding, as shown in Figure 1. In summary, our contributions in this work include:
@@ 2 Related Work
Legal Benchmarks. Recently, a series of legal benchmarks have emerged (see Table 1). They have made significant contributions to evaluating LLMs’ performance kanapala2019text; yao2025elevating, including retrieval (STARD, LeCaRD) su2024stard; li2024lecardv2, question answering (JEC-QA, Legal CQA) zhong2020jec; askari2022expert, classification (LexGLUE) chalkidis2021lexglue, reasoning (LegalBench, LexEval) guha2023legalbench; li2024lexeval, and others (Laiw and LawBench) dai2025laiw; fei2024lawbench. However, most benchmarks remain task-oriented and outcome-focused, offering limited insight into the underlying legal general intelligence of LLMs  yue2024lawllm.
Expert-level Benchmarks. To usher in the second half of AI, a series of expert-level benchmarks for evaluating LLMs have emerged across various domains cao2025toward; ni2025survey; li2025legalagentbench; li2025fundamental: PhysBench chow2025physbench and PhysReason zhang2025physreason enhance LLMs’ understanding of physics; MedXpertQA zuo2025medxpertqa and Medagentsbenchtang2025medagentsbench focus on medical knowledge; UGMathBench xu2025ugmathbench assesses math reasoning; and UniToMBench thiyagarajan2025unitombench improves theory of mind. Benchmarks like ShotBench liu2025shotbench, FinTMMBench zhu2025fintmmbench, and Chengyu-Bench fu2025chengyu evaluate other fields. However, an expert-level benchmark for legal intelligence is absent.  wang2025survey.
@@ 3 LexGenius Framework
3.1 Dimension: Education and Career Focus
The seven legal dimensions of LexGenius are based on Bloom’s Taxonomy of Educational Objectives bloom1956taxonomy, covering the cognitive hierarchy of remembering, understanding, applying, analyzing, evaluating, and creating, alongside the modular model used in legal evaluations across countries, focusing on normative understanding, rule application, procedural operation, and value judgment wu2012regulatory; moon2020case; parsons2024georgia. In the hierarchy, remembering and understanding correspond to legal understanding, applying to legal application, analyzing to legal reasoning, evaluating to legal ethics and law and society, creating to advanced arguments, legal language to clarity, and judicial practice to procedural integrity, forming a framework aligned with cognitive principles and professional needs.
@@ 3.1 Dimension: Education and Career Focus
3.2 Task: Theory and Practice Synergy
Further, based on Legal Hermeneutics leyh2021legal and Problem-Solving Cycle theory stein1993ideal, we decompose LexGenius’s 7 legal intelligence dimensions into 11 tasks. These tasks align with common legal practice requirements and focus on textual deconstruction, case adaptation, and procedural implementation. Legal Hermeneutics guides the understanding of provisions, critical analysis of texts, and terminology, ensuring accuracy in interpretation. Problem-Solving Cycle theory simulates legal practice, driving reasoning and application analysis for problem-solving, legal and ethical judgment, moral boundaries for value calibration, case reasoning, judicial procedure understanding for validation, and legal impact and social change review, forming a task system for legal problem-solving.
@@ 3.3 Ability: Constructivist Learning-based
Furthermore, based on Constructivist Learning Theory ariati2025constructivist, we extract twenty atomic legal intelligence abilities from the eleven tasks. The theory shifts from outcome assessment to capturing knowledge paths through cognitive traces, simulating real legal scenarios to ensure that the evaluation reflects true professional abilities while aligning cognitive principles with occupational needs. The hierarchy from dimensions to abilities allows LexGenius to perform a detailed, multi-dimensional assessment of LLMs’ legal intelligence, supporting evaluation and optimization.
@@ 4.1 Construction Principles
To avoid data leakage and contamination, LexGenius was built from scratch, using recent Chinese legal exam questions and judgment cases. We avoided reusing existing legal datasets to minimize contamination risks and ensure originality. A structured legal competency framework, developed by experienced legal experts, was used to comprehensively cover core legal intelligence abilities li2025survey. To ensure benchmark effectiveness, a structured review process was established mohammadi2025evaluation; li2025generation. Reviewers are master’s candidates in law, systematically trained and thoroughly familiar with key regulations, case frameworks, and legal reasoning methods.
@@ 4.2 Construction Workflow
Step 2: Construction of MCQs. There are two methods for constructing multiple-choice questions, both based on the large language model. One is to screen and modify legal examination questions. Questions meeting the legal abilities are retained; others are modified using LLM. For questions related to legal soft intelligence, an LLM is used to generate them. The first method selects and modifies questions from the legal question bank: MCQs with a single correct answer are retained. In contrast, those with multiple answers are adapted using LLMs and prompt templates to ensure fairness and difficulty. For LLM-generated MCQs, we designed prompt templates with task constraints, ability descriptions, and examples to guide question generation based on specific legal cases, ensuring unique answers and a clear legal basis.
@@ 4.4 Evaluation Method
To evaluate the LLM’s legal intelligence, LexGenius adopts a three-level framework kahan2015laws; huang2024cmdl. The dimension level categorizes tasks into legal cognition areas, providing insight into performance. The task level breaks dimensions into real-world tasks, testing the application, while the ability level evaluates legal abilities, identifying performance differences. At the ability level, scores $A_{i,j,k}$ are the average correctness of MCQs in each ability, where $A_{i,j,k}=\frac{1}{n_{i,j,k}}\sum_{m=1}^{n_{i,j,k}}C_{i,j,k,m}$, with $C_{i,j,k,m}$ the correctness of the $m$-th MCQ in the $k$-th ability of the $j$-th task in the $i$-th dimension, and $n_{i,j,k}$ the number of MCQs for that ability. At the task level, the task score $T_{i,j}$ is the average of ability scores within the task, calculated as $T_{i,j}=\frac{1}{m_{i,j}}\sum_{k=1}^{m_{i,j}}A_{i,j,k}$, where $m_{i,j}$ is the number of abilities in the $j$-th task of the $i$-th dimension. At the dimension level, the dimension score $D_{i}$ is the average of task scores, expressed as $D_{i}=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}T_{i,j}$, where $n_{i}$ is the number of tasks in the $i$-th dimension.
@@ 5.1 Experimental Setup
We evaluated twelve SOTA LLMs with LexGenius, which include DeepSeek-LLM-7B-Chat (DeepSeek-7B) bi2024deepseek, Qwen-2.5-7B-Instruct (Qwen-2.5-7B) hui2024qwen2, Qwen-2.5-1.5B-Instruct (Qwen-2.5-1.5B) hui2024qwen2, Qwen-3-8B yang2025qwen3, Qwen-3-4B yang2025qwen3, GLM-4-9B-Chat (GLM-4-9B) glm2024chatglm, LLaMA-3.2-1B-Instruct (LLaMA-3.2-1B) grattafiori2024llama, LLaMA-3.2-8B-Instruct (LLaMA-3.2-8B) grattafiori2024llama, DeepSeek-R1 guo2025deepseek, DeepSeek-V3 liu2024deepseek, GPT-4o mini hurst2024gpt, and GPT‑4.1 nano brown2020language. We followed the official protocols, using official APIs or LLM weights where applicable. The evaluation utilized two types of prompts: the first type was the Naive prompt; the second was the CoT prompt, encouraging the LLMs to perform step-by-step reasoning. To prevent potential bias, we shuffled the answer options twice and averaged the scores of each LLM.
@@ 5.5 With Different Enhanced Methods (RQ4)
As shown in Table 5, the comparison results (more details are in Appendix F) reveal a triple decoupling phenomenon in LLM legal intelligence: (i) Scale-performance decoupling shows that the Qwen2.5 series outperforms the Qwen3 series, with legal intelligence exhibiting a non-linear relationship with model size, indicating that domain-specific pretraining, not scale, drives legal abilities. (ii) Reasoning paradigm decoupling shows that CoT leads to negative transfer in legal tasks due to deterministic constraints, causing probabilistic exploration and semantic drift in closed-solution tasks. (iii) Optimization strategy decoupling shows that supervised fine-tuning benefits weak LLMs but causes catastrophic forgetting in strong ones, while the RAG method reveals the orthogonality of knowledge retrieval and reasoning. Only GRPO (reinforcement learning) achieves stable improvement by aligning reward and evaluative capacity.
@@ 6 Conclusion
Limitations
In Appendix H, we discuss the limitations of LexGenius. Furthermore, we outline the future work of LexGenius in Appendix I.
Ethical Considerations
This work complies with the ACL Ethics Policy, relying on anonymized, publicly available legal resources to ensure privacy and academic integrity.
@@ A.1 Design motivation
Legal general intelligence is not a stack of tasks, but a simulation of cognitive collaborative chains. Existing frameworks often focus on classification or question-answering tasks, presenting only macro-level accuracy on isolated benchmarks. Such metrics fail to pinpoint errors within the complex cognitive chain of legal decision-making. Legal intelligence is not a sum of discrete tasks but an organic coordination of systematic capabilities, spanning statutory interpretation, fact extraction, rule adaptation, ethical judgment, and social impact assessment. LLM performance should therefore be decomposed into a multi-stage flow, rather than treated as monolithic. Accordingly, effective evaluation must map this cognitive chain, avoiding the pitfall of reducing legal intelligence to task-solving while ignoring how the model thinks.
@@
From performance reporting to capability diagnosis and explanation. Traditional metrics like accuracy or F1 indicate output correctness, but they fail to address a fundamental question: At which cognitive stage did the model fail? Was it a semantic misunderstanding? Rule misapplication? Or an ethical blind spot? LexGenius introduces 20 atomic legal general intelligence abilities and establishes an interpretable, traceable, and auditable evaluation system through a tri-level mapping mechanism across the ability layer, task layer, and cognitive dimension. This structure reveals deficiencies in specific micro-level abilities and provides an actionable feedback loop for capability-oriented training, prompt optimization, and safety enhancement.
@@
Enabling cross-stage cognitive analysis and transferability research. The complexity of legal reasoning lies in its chained structure: statutory semantics $\rightarrow$ case fact abstraction $\rightarrow$ rule application $\rightarrow$ ethical judgment $\rightarrow$ precedent alignment. Frameworks failing to distinguish performance across these stages cannot support research in multi-hop reasoning, chain-of-thought attention, or multi-task learning. Through hierarchical abstraction and modular decomposition, our framework standardizes this cognitive pathway, offering a clear baseline for investigating knowledge transfer and generalizable reasoning capabilities.
@@
Toward a professional-grade legal general intelligence evaluation paradigm. As LLMs approach the professional thresholds of bar examinations and real-world legal practice, evaluation frameworks must likewise advance to a professional-grade level. LexGenius draws on standards from the National Legal Professional Qualification Examination and international bar exams, distinguishing specialized dimensions of legal competence, such as legal semantic understanding, norm alignment, ethical judgment, and procedural compliance. This framework breaks from the general-purpose perspective of traditional NLP benchmarks, aligning with the practical demands of legal work. Such a professional, ability-oriented, structured evaluation reveals current model boundaries and illuminates potential development trajectories.